CEPS Group is going to pay an annual dividend of OMR 2.27 a share on its common stock next year. This year, the company paid a dividend of OMR 2.15 a share. The company adheres to a constant rate of growth dividend policy. What will one share of this common stock be worth five years from now if the applicable discount rate is 8.75 percent?
Answer #1

Growth Rate = (2.27-2.15)/2.15 = 0.05581

Dividend 6 years from now = This Year's Dividend*(1.05581^6) = 2.15*1.3852 = 2.978

Share Price 5 Years from now = Dividend 6 years from now/(Discount Rate-Growth Rate) = 2.978/(0.0875-0.05581) = $93.97(approximately)